# Yibo Shi, Braghadeesh Lakshminarayanan and Cristian R. Rojas

### v1

@@ Abstract
Inference and estimation are fundamental aspects of statistics, system identification and machine learning. For most inference problems, prior knowledge is available on the system to be modeled, and Bayesian analysis is a natural framework to impose such prior information in the form of a prior distribution. However, in many situations, coming out with a fully specified prior distribution is not easy, as prior knowledge might be too vague, so practitioners prefer to use a prior distribution that is as ‘ignorant’ or ‘uninformative’ as possible, in the sense of not imposing subjective beliefs, while still supporting reliable statistical analysis. Jeffreys prior is an appealing uninformative prior because it offers two important benefits: (i) it is invariant under any re-parameterization of the model, (ii) it encodes the intrinsic geometric structure of the parameter space through the Fisher information matrix, which in turn enhances the diversity of parameter samples. Despite these benefits, drawing samples from Jeffreys prior is a challenging task. In this paper, we propose a general sampling scheme using the Metropolis-Adjusted Langevin Algorithm that enables sampling of parameter values from Jeffreys prior, and provide numerical illustrations of our approach through several examples.
@@ I Introduction
Mathematical models are essential tools for analyzing, predicting, and controlling complex physical processes. System identification is a discipline that deals with the construction of such models from experimental input-output data [1, 2]. A central task within system identification is parameter estimation: given a chosen (or presumed) model structure, the goal is to infer its unknown parameters from observed data. Classical approaches include Prediction Error Methods [1], Instrumental Variables [3], Subspace Methods [4], and Maximum Likelihood Estimation [5]. These methods typically yield point estimates that are often consistent and sometimes asymptotically efficient. Nonetheless, because the dataset used for estimation is inherently finite, there remains a level of uncertainty in the resulting parameter estimates that must be addressed. To rigorously capture and incorporate this uncertainty, Bayesian parameter estimation provides a systematic probabilistic framework, enhancing inference reliability and robustness.
@@
Among all possible prior probability densities, Jeffreys prior [9, 10] offers several advantages: (i) it has minimal subjective influence, making it a suitable default in the absence of strong domain-specific information; (ii) its form remains consistent under invertible transformations of the parameter, thus, in practice, Jeffreys prior ensures that posterior updates remain invariant to arbitrary parameter transformations, avoiding the bias introduced by coordinate choices; and (iii) the prior is proportional to the square root of the determinant of the Fisher Information Matrix (FIM), thus it encodes the intrinsic geometric structure of the parameter space. This last property is particularly valuable in data-driven parameter estimation [11], where the induced Riemannian geometry facilitates sampling diverse parameter values. Such diversity enhances the robustness of machine learning models trained on these samples, as the parameters reflect distinct regions. Despite these advantages, Jeffreys prior faces practical limitations. Computing the Fisher information matrix (FIM) in closed form may be intractable or prohibitively expensive for complex or high-dimensional models. While numerical approximations can yield partial FIM estimates, they introduce additional uncertainty in the sampling process: the derived potential function is subject to estimation errors, potentially degrading performance in naive sampling schemes. Furthermore, Jeffreys prior’s normalizing constant is rarely known explicitly, even if the parameter space is bounded and the prior is integrable, complicating direct sampling.
@@
These difficulties are typically addressed through Markov chain Monte Carlo (MCMC) methods, which are widely used for sampling from challenging or high-dimensional distributions [12]. Among these, Langevin-based Monte Carlo (LMC) [13] employs gradient information to guide proposals efficiently in parameter space. Most research on advanced variants addresses some specific challenges of LMC. For instance, Constrained Ensemble LMC [14] ensures physically or statistically valid sample updates by projecting the samples onto constraint sets at each iteration; the Metropolis–Adjusted Langevin Algorithm (MALA) augments LMC with a Metropolis–Hastings accept/reject step, correcting discretization errors and guaranteeing exact convergence to the target distribution regardless of moderate step-size mis-specifications [13, 15]. Among the LMC variants, MALA offers a straightforward way to handle constraints and correct for discretization errors while preserving the geometric advantages of Jeffreys prior via a gradient-based proposal mechanism.
@@
In this paper, we introduce a MALA-based scheme for sampling Jeffreys prior in two distinct scenarios: (i) when the FIM can be derived analytically, and (ii) when an analytical form of the FIM is unavailable, but the score function can be estimated via particle filtering [16, 17], allowing us to compute an approximate FIM that is then used within MALA. Furthermore, we demonstrate an application of Jeffreys prior to generate a diverse set of parameter samples to enhance the performance of a data-driven estimator, by providing improved estimates of the parameters of a model. Our main contributions include:
@@
  • Extension of this scheme to nonlinear dynamical systems through particle-filter-based FIM estimation;
@@
The remainder of the paper is organized as follows: in Section II, we define the problem statement. Section III introduces the proposed algorithm to sample from Jeffreys prior based on MALA, while Section IV provides several numerical illustrations of our method. Finally, we conclude the paper in Section V.
@@ II Problem statement
where $y\in\mathcal{Y}$ denotes the observations, $\nabla_{\boldsymbol{\theta}}$ denotes the gradient with respect to $\boldsymbol{\theta}$, and the expectation operator $\mathbb{E}_{y\sim p(\cdot;\,\boldsymbol{\theta})}[\cdot]$ is defined as
[…]
@@
where $\boldsymbol{J}_{\boldsymbol{\theta}}$ is the FIM defined in (1) and $\det(\cdot)$ denotes the determinant function.
@@
Sampling from Jeffreys prior is thus highly relevant in system identification, machine learning, and statistics, but poses two main challenges (i) for complex models, computing or approximating the FIM can introduce uncertainty, and (ii) the normalizing constant of Jeffreys prior is rarely known, complicating direct sampling. The objective of this paper is to address these challenges by proposing an LMC approach that generates Markovian samples whose stationary distribution corresponds precisely to Jeffreys prior $\pi(\boldsymbol{\theta})$.
@@ III Proposed Method
In this section, we present the MALA-based method for sampling from Jeffreys prior. The proposed approach accommodates two scenarios: one where the FIM can be computed analytically, and the other where it must be approximated. By unifying these cases under the MALA framework, we enable sampling from the Jeffreys prior across a broad class of parameterized systems, regardless of whether the FIM has a closed-form or is numerically derived.
@@ III-A Langevin-based Monte Carlo
To verify that the SDE in (4) samples from the Jeffreys prior defined in (3), we use the corresponding Fokker–Planck equation [20]. For simplicity, we first consider the one-dimensional case with parameter $\theta$. The Fokker–Planck equation associated with the Langevin SDE
@@
describes the time evolution of the probability density $p(\theta,t)$ according to [20]
@@
In the steady-state regime $\partial p(\theta,t)/\partial t=0$, the probability density reaches a stationary distribution denoted by $p_{\infty}(\theta)$. Imposing zero probability flux at the boundaries (or as $|\theta|\to\infty$), we obtain
$p_{\infty}(\theta)\;\propto\;\exp\bigl{(}-V(\theta)\bigr{)}.$
Hence, the normalized stationary distribution is
$p_{\infty}(\theta)\;=\;\frac{\exp\bigl{(}-V(\theta)\bigr{)}}{Z},$
@@
$V(\theta)=-\frac{1}{2}\ln\det(J_{\theta}),$ (7)
then it follows from (III-A) that
$p_{\infty}(\theta)\;=\;\frac{1}{Z}\sqrt{\det(J_{\theta})}\;\propto\;\pi(\theta).$
Thus, the stationary distribution associated with the Langevin SDE (4) exactly coincides with Jeffreys prior. This result is a classical consequence of the relationship between the Fokker–Planck equation and its associated Itô diffusion [22], which underlies the theoretical foundations of LMC methods [13, 23].
Using the update rule in (6) directly to sample from $\pi(\boldsymbol{\theta})$ is referred to as Unadjusted Langevin Algorithm (ULA). It performs an iteration of (6) to generate the samples $\boldsymbol{\theta}_{i}$ for $i=0,1,2,\dots$, which approximates the target distribution $\pi(\boldsymbol{\theta})$ after a sufficient burn-in period. However, ULA is highly sensitive to the choice of step size $\tau$. If $\tau$ is excessively large, discretization errors introduce bias into the invariant distribution or even cause divergence; if too small, the chain suffers slow mixing, resulting in computational inefficiency. Moreover, practical applications often require the sampled parameter $\boldsymbol{\theta}_{i}$ to stay in the constrained parameter space $\Theta_{c}\subset\Theta$ due to physical limitations or prior knowledge. ULA lacks a mechanism to enforce these constraints, potentially proposing values outside the feasible region. In the following subsection, we describe the MALA framework, which addresses these limitations, providing robust and efficient sampling from Jeffreys prior.
@@ III-B Metropolis-Adjusted Langevin Algorithm
MALA improves upon ULA by introducing a Metropolis–Hastings accept/reject step that compensates for the errors induced by the discretization. Note that in (6), the proposed variable $\boldsymbol{\theta}_{t+1}$ follows the Gaussian distribution:
$\boldsymbol{\theta}_{t+1}\sim\mathcal{N}(\boldsymbol{\theta}_{t}-\tau\nabla_{\boldsymbol{\theta}}V(\boldsymbol{\theta}_{t}),2\tau).$ (8)
@@
$q\bigl{(}\boldsymbol{\theta}_{t+1}\mid\boldsymbol{\theta}_{t}\bigr{)}\propto\exp\!\Bigl{(}-\frac{\|\,\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}+\tau\nabla_{\boldsymbol{\theta}}V(\boldsymbol{\theta}_{t})\|_{2}^{2}}{4\,\tau}\Bigr{)}.$ (9)
Given a proposed sample $\boldsymbol{\theta}^{\prime}$, MALA accepts it with probability [15]
$\rho^{\text{MALA}}(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{t})\;=\;\min\!\Bigl{\{}1,\;\frac{\exp(-V(\boldsymbol{\theta}^{\prime}))\,q\bigl{(}\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}^{\prime}\bigr{)}}{\exp(-V(\boldsymbol{\theta}_{t}))\,q\bigl{(}\boldsymbol{\theta}^{\prime}\mid\boldsymbol{\theta}_{t}\bigr{)}}\Bigr{\}}.$ (10)
If the proposal is accepted, the chain advances as $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}^{\prime}$; otherwise, it remains at the current position, $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}$. This ensures that the chain remains exactly invariant with respect to $\pi(\boldsymbol{\theta})$ under mild conditions, yielding more robust sampling compared to ULA [15]. Furthermore, MALA inherently accommodates constrained parameter spaces $\boldsymbol{\theta}\in\Theta_{\text{c}}$ by modifying the acceptance probability as follows:
$\rho_{c}^{\text{MALA}}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{t}\right)=\begin{cases}\rho^{\text{MALA}}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{t}\right),&\text{if }\boldsymbol{\theta}^{\prime}\in\Theta_{c},\\ 0,&\text{otherwise }.\end{cases}$
Therefore, the accept/reject step based on $\rho_{c}^{\text{MALA}}$ automatically discards proposals that violate these bounds, thereby preserving both the feasibility and the accuracy of the sampling process.
@@ III-C Sampling with Estimated FIM
For certain systems with an analytical form of the FIM, we can directly define our potential function as the multivariate form as a generalization of (7),
@@
with its gradient as
@@
In many dynamic systems, particularly nonlinear state-space (NLSS) models, closed-form expressions for FIMs are not available. In such cases, one must resort to numerical approximations. In this paper, we adopt a particle-filter–based approach to estimate the FIM, following [17], whereby Forward Filtering–Backward Smoothing (FFBSm) provides an unbiased Monte Carlo approximation of the score function, and consequently, the FIM. Despite the inherent stochasticity of particle filters, the resulting estimates are consistent and converge to the true FIM under standard regularity conditions as the number of particles increases.
Remark 1. For a concrete illustration of particle-filter–based FIM estimation within our framework, we refer the reader to the NLSS model simulation in Section IV.
Given the estimated FIM $\boldsymbol{\hat{J}}{\boldsymbol{\theta}}$, the gradient ${\partial\boldsymbol{J}_{\boldsymbol{\theta}}}/{\partial\boldsymbol{\theta}}$ can be approximated by the one-point unbiased estimate. Specifically, we introduce a random perturbation $\boldsymbol{\mu}\sim\mathcal{N}(\boldsymbol{0}_{d},\boldsymbol{I}_{d})$ and approximate the derivative as
@@
where $\delta>0$ is a small step size, and $\boldsymbol{\hat{J}}_{\boldsymbol{\theta}+\delta\boldsymbol{\mu}}$ is computed using the same estimation procedure as for $\boldsymbol{\hat{J}}_{\boldsymbol{\theta}}$. This estimator is unbiased in expectation (for $\delta\to 0$) and substantially reduces the computational burden compared to coordinate-wise finite differences.
Remark 2. Although the use of such an approximate gradient in (13) might introduce errors, the inclusion of a Metropolis–Hastings accept/reject step in our MALA-based scheme compensates for these inaccuracies and ensures that the target distribution $\pi(\boldsymbol{\theta})$ is still preserved [24]. Standard results in MCMC theory (see, e.g., [13, 15]) guarantee that, as long as the gradient estimator is unbiased and its noise is properly accounted for by the acceptance step, the resulting Markov chain will have $\pi(\boldsymbol{\theta})$ as its stationary distribution. This ensures that our methodology accommodates both closed-form and numerically approximated FIMs without sacrificing correctness.
Algorithm 1 integrates the particle-filter–based FIM estimation with the one-point gradient approximation within MALA. In this way, even when the FIM is only available approximately via particle methods, our approach guarantees convergence to Jeffreys prior. This enables robust sampling for complex nonlinear state-space models, thereby enhancing both prediction and inference in system identification.
@@
1:Initial guess $\boldsymbol{\theta}_{0}$, possible constrained parameter space $\Theta_{\text{c}}$, step size $\tau$, number of iterations $N$, small finite-difference parameter $\delta>0$

### v3

@@ Abstract
Inference and estimation are fundamental in statistics, system identification, and machine learning. When prior knowledge about the system is available, Bayesian analysis provides a natural framework for encoding it through a prior distribution. In practice, such knowledge is often too vague to specify a full prior distribution, motivating the use of default “uninformative” priors that minimize subjective bias. Jeffreys prior is an appealing uninformative prior because: (i) it is invariant under any re-parameterization of the model, (ii) it encodes the intrinsic geometric structure of the parameter space through the Fisher information matrix, which in turn enhances the diversity of parameter samples. Despite these benefits, drawing samples from Jeffreys prior is challenging. In this paper, we develop a general sampling scheme using the Metropolis-Adjusted Langevin Algorithm that enables sampling of parameter values from Jeffreys prior; the method extends naturally to nonlinear state–space models. The resulting samples can be directly used in sampling-based system identification methods and Bayesian experiment design, providing an objective, information-geometric description of parameter uncertainty. Several numerical examples demonstrate the efficiency and accuracy of the proposed scheme.
@@ I Introduction
Mathematical models are essential tools for analyzing, predicting, and controlling complex physical processes. System identification is a discipline that deals with the construction of such models from experimental input-output data [1, 2]. A central task within system identification is parameter estimation: given a chosen (or presumed) model structure, the goal is to infer its unknown parameters from observed data. Classical approaches include Prediction Error Methods [1], Instrumental Variables [3], Subspace Methods [4], and Maximum Likelihood Estimation [5]. These methods typically yield point estimates that are often consistent and even asymptotically efficient. However, as the dataset used for estimation is inherently finite, there remains a level of uncertainty in the resulting parameter estimates that must be addressed. To rigorously capture this uncertainty, Bayesian parameter estimation provides a systematic probabilistic framework, enhancing inference reliability and robustness.
@@
Among all possible prior probability densities, Jeffreys prior [9, 10] offers several advantages: (i) it has minimal subjective influence, making it a suitable choice in the absence of strong domain-specific information; (ii) its form remains consistent under invertible transformations of the parameter, thus it ensures that posterior updates remain invariant to arbitrary parameter transformations, avoiding the bias introduced by coordinate choices; and (iii) the prior is proportional to the square root of the determinant of the Fisher Information Matrix (FIM), which encodes the intrinsic geometric structure of the parameter space. This last property is particularly valuable in data-driven parameter estimation [11], where the induced Riemannian geometry facilitates sampling diverse parameter values. Such diversity enhances the robustness of machine learning models trained on these samples, as the parameters reflect distinct regions. Despite these advantages, Jeffreys prior faces practical limitations. Computing the FIM in closed form may be intractable or prohibitively expensive for complex or high-dimensional models. While numerical approximations can yield partial FIM estimates, they introduce additional uncertainty in the sampling process: the derived potential function is subject to estimation errors, potentially degrading performance in naive sampling schemes. Furthermore, Jeffreys prior’s normalizing constant is rarely known explicitly, even if the parameter space is bounded and the prior is integrable, complicating direct sampling.
@@
These difficulties are typically addressed through Markov chain Monte Carlo (MCMC) methods, which are widely used for sampling from challenging or high-dimensional distributions [12]. Among these, Langevin-based Monte Carlo (LMC) [13] employs gradient information to guide proposals efficiently in parameter space. Several variants have been proposed to address some specific challenges of LMC. For instance, Constrained Ensemble LMC [14] ensures physically valid sample updates by projecting the samples onto constraint sets at each iteration; the Metropolis–Adjusted Langevin Algorithm (MALA) augments LMC with a Metropolis–Hastings accept/reject step, correcting discretization errors and guaranteeing exact convergence to the target distribution regardless of moderate step-size mis-specifications [13, 15]. Among the LMC variants, MALA offers a simple way to handle constraints and correct for discretization errors while preserving the geometric advantages of Jeffreys prior via a gradient-based proposal mechanism.
@@
In this paper, we introduce a MALA-based scheme for sampling Jeffreys prior in two distinct scenarios: (i) when the FIM can be derived analytically, and (ii) when an analytical form of the FIM is unavailable, but the score function can be estimated via particle filtering (PF) [16, 17], allowing us to compute an approximate FIM that is then used within MALA. Furthermore, we demonstrate an application of Jeffreys prior to generate a diverse set of parameter samples to enhance the performance of a data-driven estimator, by providing improved estimates of the parameters of a model. Our main contributions include:
@@
  • Extension of this scheme to nonlinear dynamical systems through PF-based FIM estimation;
@@
The remainder of the paper is organized as follows: in Section II, we define the problem statement. Section III introduces the proposed algorithm to sample from Jeffreys prior, while Section IV provides several numerical illustrations of our method. Finally, we conclude the paper in Section V.
@@ II Problem statement
where $y\in\mathcal{Y}$ denotes the observations, $\nabla_{\boldsymbol{\theta}}$ denotes the gradient with respect to $\boldsymbol{\theta}$, and $\mathbb{E}_{y\sim p(\cdot;\,\boldsymbol{\theta})}[\cdot]=\int_{\mathcal{Y}}\bigl{[}\cdot\bigr{]}\,p(y;\,\boldsymbol{\theta})\,dy.$
@@
where $\det(\cdot)$ denotes the determinant function.
@@
Sampling from Jeffreys prior is thus highly relevant in system identification, machine learning, and statistics, but poses two main challenges: (i) for complex models, computing or approximating the FIM can introduce uncertainty, and (ii) the normalizing constant of Jeffreys prior is rarely known, complicating direct sampling. The objective of this paper is to address these challenges by proposing an LMC approach that generates Markovian samples whose stationary distribution corresponds precisely to Jeffreys prior $\pi(\boldsymbol{\theta})$.
@@ III Proposed Method
In this section, we first describe two standard sampling methods: LMC, and its improved variant, MALA. Then, in Subsection III-C, we introduce a novel MALA-based method for sampling from Jeffreys prior, which accommodates two scenarios: one where the FIM can be computed analytically, and the other where it must be approximated.
@@ III-A Langevin-based Monte Carlo
To verify that the SDE in (4) samples from the Jeffreys prior defined in (3), we use the corresponding Fokker–Planck equation [20]. For simplicity of illustration, we consider the one-dimensional case with parameter $\theta$. The Fokker–Planck equation [20] associated with the Langevin SDE
@@
describes the evolution of the probability density $p(\theta,t)$ as
@@
In the steady-state regime $\partial p(\theta,t)/\partial t=0$, the probability density reaches a stationary distribution denoted by $p_{\infty}(\theta)$. Imposing zero probability flux at the boundaries (or as $|\theta|\to\infty$), we obtain $p_{\infty}(\theta)\;\propto\;\exp\bigl{(}-V(\theta)\bigr{)}$. Hence, the normalized stationary distribution is
$p_{\infty}(\theta)\;=\;\frac{\exp\bigl{(}-V(\theta)\bigr{)}}{Z},$ (7)
@@
$V(\theta)=-\frac{1}{2}\ln(J_{\theta}),$ (8)
then it follows from (7) that $p_{\infty}(\theta)\;\propto\;\pi(\theta).$ Thus, the stationary distribution associated with the Langevin SDE (4) coincides exactly with Jeffreys prior.
The update rule in (6) to sample from $\pi(\boldsymbol{\theta})$ leads to the Unadjusted Langevin Algorithm (ULA), which iterates (6) to generate the samples $\boldsymbol{\theta}_{i}$ for $i=0,1,2,\dots$, whose distribution approximates $\pi(\boldsymbol{\theta})$ after a sufficient burn-in period. However, ULA is highly sensitive to the choice of step size $\tau$. If $\tau$ is excessively large, discretization errors introduce bias into the invariant distribution or even cause divergence; if too small, the chain suffers slow mixing, resulting in computational inefficiency. Moreover, one often requires the sampled parameter $\boldsymbol{\theta}_{i}$ to stay in a constrained parameter space $\Theta_{c}\subset\Theta$ due to physical limitations or prior knowledge. ULA lacks a mechanism to enforce these constraints. In the following subsection, we describe the MALA framework, which addresses these limitations.
@@ III-B Metropolis-Adjusted Langevin Algorithm
MALA [15] improves upon ULA by introducing a Metropolis–Hastings accept/reject step that compensates for the errors induced by the discretization. Note that in (6), the proposed variable $\boldsymbol{\theta}_{i+1}$ follows the Gaussian distribution
$\boldsymbol{\theta}_{i+1}\sim\mathcal{N}(\boldsymbol{\theta}_{i}-\tau\nabla_{\boldsymbol{\theta}}V(\boldsymbol{\theta}_{i}),2\tau).$ (9)
@@
$q\bigl{(}\boldsymbol{\theta}_{i+1}\mid\boldsymbol{\theta}_{i}\bigr{)}\propto\exp\Bigl{(}-\frac{\|\,\boldsymbol{\theta}_{i+1}-\boldsymbol{\theta}_{i}+\tau\nabla_{\boldsymbol{\theta}}V(\boldsymbol{\theta}_{i})\|_{2}^{2}}{4\,\tau}\Bigr{)}.$ (10)
MALA accepts a given proposed sample $\boldsymbol{\theta}^{\prime}$ with probability
$\rho^{\text{MALA}}(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{i})\;=\;\min\Bigl{\{}1,\;\frac{\exp(-V(\boldsymbol{\theta}^{\prime}))\,q\bigl{(}\boldsymbol{\theta}_{i}\mid\boldsymbol{\theta}^{\prime}\bigr{)}}{\exp(-V(\boldsymbol{\theta}_{i}))\,q\bigl{(}\boldsymbol{\theta}^{\prime}\mid\boldsymbol{\theta}_{i}\bigr{)}}\Bigr{\}}.$
If the proposal is accepted, the chain advances as $\boldsymbol{\theta}_{i+1}=\boldsymbol{\theta}^{\prime}$; otherwise, it remains at the current position, $\boldsymbol{\theta}_{i+1}=\boldsymbol{\theta}_{i}$. This ensures that the chain remains exactly invariant with respect to $\pi(\boldsymbol{\theta})$ under mild conditions, yielding more robust sampling compared to ULA [15]. Furthermore, MALA inherently accommodates constrained parameter spaces $\boldsymbol{\theta}\in\Theta_{\text{c}}$ by modifying the acceptance probability as follows:
$\rho_{c}^{\text{MALA}}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{i}\right)=\begin{cases}\rho^{\text{MALA}}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}_{i}\right),&\text{if }\boldsymbol{\theta}^{\prime}\in\Theta_{c},\\ 0,&\text{otherwise }.\end{cases}$
@@ III-C Sampling with Estimated FIM
If an analytical form of the FIM is available, we can directly define the multivariate potential function
@@
which is a generalization of (8), with gradient
@@
For many problems, such as nonlinear state-space (NLSS) models, closed-form expressions for FIMs are not available. In such cases, one must resort to numerical approximations. In this paper, we adopt a PF–based approach to estimate the FIM, following [17], whereby Forward Filtering–Backward Smoothing (FFBSm) provides an unbiased Monte Carlo approximation of the score function, and consequently, the FIM. Despite the inherent stochasticity of particle filters, the resulting estimates are consistent and converge to the true FIM under standard regularity conditions as the number of particles increases.
Remark III.1
For a concrete illustration of PF–based FIM estimation within our framework, we refer the reader to the NLSS model simulation in Section IV.
Given the estimated FIM $\boldsymbol{\hat{J}}_{\boldsymbol{\theta}}$, the gradient ${\partial\boldsymbol{J}_{\boldsymbol{\theta}}}/{\partial\boldsymbol{\theta}}$ can be approximated by the one-point unbiased estimate. Specifically, we introduce a random perturbation $\boldsymbol{\mu}\sim\mathcal{N}(\boldsymbol{0}_{d},\boldsymbol{I}_{d})$ and approximate the derivative as
@@
where $\delta>0$ is a small step size, and $\boldsymbol{\hat{J}}_{\boldsymbol{\theta}+\delta\boldsymbol{\mu}}$ is computed using the same estimation procedure as for $\boldsymbol{\hat{J}}_{\boldsymbol{\theta}}$. This estimator is unbiased (for $\delta\to 0$) and substantially faster compared to coordinate-wise finite differences.
Remark III.2
Although the use of the approximate gradient in (13) might introduce errors, the Metropolis–Hastings step compensates for these inaccuracies and ensures that the target distribution $\pi(\boldsymbol{\theta})$ is still preserved [22]. Standard results in MCMC theory [13, 15] guarantee that, as long as the gradient estimator is unbiased and its noise is properly accounted for by the acceptance step, the resulting Markov chain will have $\pi(\boldsymbol{\theta})$ as its stationary distribution.
Algorithm 1 integrates the PF–based FIM estimation with the one-point gradient approximation within MALA. In this way, even when the FIM is only approximately available, our approach guarantees convergence to Jeffreys prior.
@@
1:Initial guess $\boldsymbol{\theta}_{0}$, constraint set $\Theta_{\text{c}}$, step size $\tau$, number of iterations NN, small step size $\delta>0$